OpenStudy (anonymous):

find the length of the curve given by r(t)=ti+ln(sect)j+3k from t=0 to t=pi/4

OpenStudy (anonymous):

\[\int\limits_{0}^{\prod_/4}\left( t \right)+\ln(\sec)+3\]